OpenStudy (anonymous):

-(12x)to the power of -2

OpenStudy (zzr0ck3r):

-(12x)^(-2) = -1/(12x)^2 = -1/(12^2*x^2) = ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

it needs to be simplified

OpenStudy (zzr0ck3r):

there is one more step, and it simplified

OpenStudy (anonymous):

so....

OpenStudy (anonymous):

can u do the last step and then explain pls

OpenStudy (anonymous):

explain it

OpenStudy (turingtest):

you are combining two rules...

OpenStudy (zzr0ck3r):

do you see anything in there that can be simplified?

OpenStudy (turingtest):

x^(-a)=1/x^a and (xy)^a=(x^a)(y^a) between that and what @zzr0ck3r said you should be able to get it

OpenStudy (anonymous):

...

OpenStudy (zzr0ck3r):

what is 1/2^2?

OpenStudy (anonymous):

oooooooh

OpenStudy (anonymous):

ty

OpenStudy (zzr0ck3r):

show us what you got